OpenStudy (sparklestaraa):

Using your function, F(3)=2(3) + 5 explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences.

OpenStudy (sparklestaraa):

desperate help :(

OpenStudy (danjs):

you have a function y=f(x). f(x) = 2*x + 5 Want f(x) , when x=3 f(3) = 2*3 + 5 = 11

OpenStudy (sparklestaraa):

Ok i figured

OpenStudy (sparklestaraa):

Thank you so much

OpenStudy (sparklestaraa):

Using complete sentences, describe to the Martians how to find the inverse of your function @danjs

OpenStudy (mww):

to find an inverse you will change the subject of the function to the independent variable, x. After this, you swap x and y.

OpenStudy (sparklestaraa):

Cn you show me an example?

OpenStudy (mww):

Change the subject to x: y = 12x - 5 12x = y + 5 x = (y+5)/12 Swap x and y y = (x+5)/12 So your inverse of y = 12x - 5 is y = (x+5)/12